1 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

Bond order is the number, which indicates the total number of bonds present between two atoms. Bond order describes the bond strength of the molecule;

If the bond order is more, the bond strength would also be more for that molecule.

BondOrderBondstrength{\rm{Bond Order}} \propto {\rm{Bond}}\,{\rm{strength}} .

Fundamentals

Bond order calculation:

B.O=12(B.EA.B.E)Where,B.O=Bondorder.B.E=Bondingelectrons.A.B.E=Antibondingelectrons.\begin{array}{l}\\{\rm{B}}{\rm{.O}}\,{\rm{ = }}\frac{{\,{\rm{1}}}}{{\rm{2}}}{\rm{(B}}{\rm{.E - A}}{\rm{.B}}{\rm{.E)}}\\\\{\rm{Where,}}\\\\{\rm{B}}{\rm{.O}}\,{\rm{ = }}\,{\rm{Bond}}\,{\rm{order}}{\rm{.}}\\\\{\rm{B}}{\rm{.E}}\,{\rm{ = }}\,{\rm{Bonding}}\,{\rm{electrons}}{\rm{.}}\\\\{\rm{A}}{\rm{.B}}{\rm{.E}}\,{\rm{ = }}\,{\rm{Anti - bonding}}\,{\rm{electrons}}{\rm{.}}\\\end{array}

The general molecular orbital diagram is shown below.

is
molecule
Atom
Atom
Не
Не
Не2
Energy

The given molecule is He2{\rm{H}}{{\rm{e}}_{\rm{2}}} which has 4 electrons in it.

Molecular orbital diagram for the molecule He2{\rm{H}}{{\rm{e}}_{\rm{2}}} is shown below.

1s
|4k
He atom
He atom
He2 Molecule
Atom
Atom
Energy

The given molecule is He2{\rm{H}}{{\rm{e}}_{\rm{2}}} .

The number of bonding electrons in He2{\rm{H}}{{\rm{e}}_{\rm{2}}} molecule is 2

The number of anti-bonding electrons in He2{\rm{H}}{{\rm{e}}_{\rm{2}}} molecule is 2

Now,

B.O=12(B.EA.B.E)=12(22)=12(0)=0\begin{array}{l}\\{\rm{B}}{\rm{.O}}\,{\rm{ = }}\,\,\frac{{\,{\rm{1}}}}{{\rm{2}}}{\rm{(B}}{\rm{.E - A}}{\rm{.B}}{\rm{.E)}}\,\\\\\,\,\,\,\,\,\,\,\,{\rm{ = }}\,\,\frac{{\,{\rm{1}}}}{{\rm{2}}}{\rm{(2 - 2)}}\\\\\,\,\,\,\,\,\,\,{\rm{ = }}\,\,\frac{{\,{\rm{1}}}}{{\rm{2}}}{\rm{(0)}}\\\\\,\,\,\,\,\,\,\,{\rm{ = }}\,\,{\rm{0}}\\\end{array}

Therefore, the bond order for the given molecule He2{\rm{H}}{{\rm{e}}_{\rm{2}}} is zero.

Ans:

Bond order and the molecular orbital of the diagram is given below.

bond order
ls
0
0.5
is
4h
1
S
He atom
He atom
O 1.5
4h
He2 Molecule
Atom
Atom
Energy